# Ikeda type lift on ${\rm SO}(3,n+1)$

### v2

@@ 8.4. Hecke Eigenform Property
Since the finite part of $f$ is a non-zero section of the one-dimensional space $\Big({\rm Ind}^{G(\mathbb{A}_{\mathbb{f}})}_{P(\mathbb{A}_{\mathbb{f}})}|\nu|_{\mathbb{f}}^{l-\frac{n+2}{2}}\Big)^{G(\widehat{\mathbb{Z}})}$ which also preserved under the action of $\mathcal{H}_{\mathbb{f}}$. Thus, there exists $c_{h}\in\mathbb{C}^{\times}$ such that $T_{h}f=c_{h}f$. The claim follows. ∎
@@
By using [10, Lemma 10.1] and the argument for [15, Proposition 10.1], we see that
@@ 9.1. $n$ even
We give a conjectural Arthur parameter of $\Pi_{F}$: Let $\phi_{f}:\mathcal{L}\longrightarrow{\rm SL}_{2}(\mathbb{C})$ be the hypothetical Langlands parameter attached to $f$. Then ${\rm Sym}^{2}\phi_{f}:SL_{2}(\mathbb{C})\times{\rm SO}(3,\mathbb{C})$ parametrizes ${\rm Sym}^{2}\pi_{f}$. The distinguished unipotent orbit $(8a+3)$ of ${\rm SO}(8a+3,\mathbb{C})$ gives rise to a map $\phi_{u}:{\rm SL}_{2}(\mathbb{C})\longrightarrow{\rm SO}(8a+3,\mathbb{C})$. Then consider
@@ 10. Ikeda type lift as CAP representation
When $n=2$, then $G={\rm SO}(3,3)$ and our method (in Section 6) breaks down, and we do not get a cusp form on ${\rm SO}(3,3)$. For example, it is no longer true that ${\rm SO}(3,3)$ has a relative root system of type $B_{3}$ in Section 6.1. Thus, we cannot apply the results in Section 8.2. It agrees with the fact that ${\rm SL}_{4}$ is isogenous to ${\rm SO}(3,3)$, and ${\rm SL}_{4}$ does not have CAP representations, since any cuspidal representation of ${\rm SL}_{n}$ can be extended to a cuspidal representation of ${\rm GL}_{n}$ (cf. [3, p.236, line 9-10]). On the other hand, H. Narita, A. Pitale and S. Wagh [21] constructed a CAP representation on ${\rm O}(1,5)$ from Maass cusp form using the theta lift.
@@ 11.2.2. The case when $p=2$ (and $n$ is odd)
where the final identity uses a well-known formula for quadratic Gauss sums (cf. [2, Theorem 1.5.4, p.27]). ∎
We also need the following [2, p. 29]:
@@ 12. Appendix
In this section, we explicitly compute $C(S,r)=C(r,\sqrt{\beta}^{-1}\lambda)$ in (7.11) as a generalization of [24, Theorem A.1]: Let
@@
Let $I_{v}=I_{v}(r,\mu,\lambda)$. Then by the change of variables, we see that $I_{-v}=I_{v}$ and $\overline{I_{v}}=I_{v}$. So $I_{v}$ is real. Also [24, Proposition A.2] says, for any $r$,
@@
Now we explain how to modify the proof of [24, Proposition A.2] to obtain $I_{0}(2r,\mu,\lambda)=C(2r,\lambda)e^{-\mu}.$ Then together with the recurrence relation above, we obtain the theorem.
Let $s=\zeta t^{\frac{1}{2}}+\zeta^{-1}t^{-\frac{1}{2}}$, where $\zeta=e^{\sqrt{-1}\frac{\pi}{4}}$. Then $s^{2}=2+\sqrt{-1}(t-t^{-1})$ and $s\bar{s}=t+t^{-1}$. Let $\alpha=\mu+\lambda^{2}$. As in the proof of [24, Proposition A.2], we have

### v3

@@ 8.4. Hecke Eigenform Property
The finite part of the section $f$ is a non-zero section of the one-dimensional space $\Big({\rm Ind}^{G(\mathbb{A}_{\mathbb{f}})}_{P(\mathbb{A}_{\mathbb{f}})}|\nu|_{\mathbb{f}}^{l-\frac{n+2}{2}}\Big)^{G(\widehat{\mathbb{Z}})}$ which is also preserved under the action of $\mathcal{H}_{\mathbb{f}}$. Thus, there exists $c_{h}\in\mathbb{C}^{\times}$ such that $T_{h}f=c_{h}f$. The claim follows. ∎
@@
By using [14, Lemma 10.1] and the argument for [19, Proposition 10.1], we see that
@@
8.5. Infinity type
Let $\Pi_{F_{f}}=\Pi_{\infty}\otimes\otimes^{\prime}_{p}\Pi_{p}$ be the cuspidal automorphic representation associated to $F_{f}$. We check the claim for $\Pi_{\infty}$ in the main theorems. By assumption $l>n+1$ and Proposition 3.2, we see that $\Pi_{\infty}\simeq\Pi^{3,n+1}_{l,0}$. It follows from [22, Section 2.2] that as a $(\mathfrak{g},K)$-module, $\Pi^{3,n+1}_{l,0}$ is identified with the derived functor module $A_{\mathfrak{q}}(\lambda-\rho)=A_{\mathfrak{q}}(l-n-1)$ with $\lambda=l-\frac{n}{2}$ and $\rho=\frac{n}{2}+1$ in the notation therein. Here $\mathfrak{q}=\mathfrak{l}_{\mathbb{C}}+\mathfrak{u}$ is a $\theta$-stable parabolic subalgebra of $\mathfrak{g}_{\mathbb{C}}$ with the Levi subalgebra $\mathfrak{l}_{C}$ and the normalizer of $\mathfrak{u}$ in $G(\mathbb{R})$ is isomorphic to ${\rm SO}(2)\times{\rm O}(1,n+1)$. Thus, $\Pi_{\infty}$ is cohomological. Further, one can check $\mathfrak{l}_{\mathbb{C}},\mathfrak{l}_{\mathbb{C}}]\not\subset\mathfrak{k}:={\rm Lie}(K_{\mathbb{C}})$ by using the description of $\mathfrak{u}$ given in [22, p.7, line 8 from Section 2.2]. Thus, as explained in [33, p.58, line -9], $A_{\mathfrak{q}}(l-n-1)$ is non-tempered.
Proof of Corollary 1.4. Keep the notation as above. This kind of claims have already appeared in the proof of [24, p.180, Corollary 1.3]. It is easy to see that ${\rm dim}_{\mathbb{C}}(\mathfrak{u}\cap\mathfrak{p})=n+1$ where $\mathfrak{g}:={\rm Lie}(G(\mathbb{R}))_{\mathbb{C}}=\mathfrak{k}+\mathfrak{p}$ is the Cartan decomposition. Applying [33, p.64, Theorem 3.3] with $R={\rm dim}_{\mathbb{C}}(\mathfrak{u}\cap\mathfrak{p})=n+1$, we have
$H^{n+2}(\mathfrak{g},K,A_{\mathfrak{q}}(\lambda-\rho))\neq 0.$
Notice $\Pi_{\infty}\otimes E^{\vee}_{l}$ contains the trivial representation as a $(\mathfrak{g},K)$-module. Thus, $\Pi_{F_{f}}$ contributes to $H^{n+2}(\Gamma\backslash X,\mathcal{E}_{l})$. The claim now follows from the cuspidality of $\Pi_{F_{f}}$. $\square$
@@ 9.1. $n$ even
We give a conjectural Arthur parameter of $\Pi_{F}$: Let $\phi_{f}:\mathcal{L}\longrightarrow{\rm SL}_{2}(\mathbb{C})$ be the hypothetical Langlands parameter attached to $f$. Then ${\rm Sym}^{2}\phi_{f}:\mathcal{L}\longrightarrow{\rm SO}(3,\mathbb{C})$ parametrizes ${\rm Sym}^{2}\pi_{f}$. The distinguished unipotent orbit $(8a+3)$ of ${\rm SO}(8a+3,\mathbb{C})$ gives rise to a map $\phi_{u}:{\rm SL}_{2}(\mathbb{C})\longrightarrow{\rm SO}(8a+3,\mathbb{C})$. Then consider
@@ 10. Ikeda type lift as CAP representation
When $n=2$, then $G={\rm SO}(3,3)$ and our method (in Section 6) breaks down, and we do not get a cusp form on ${\rm SO}(3,3)$. For example, it is no longer true that ${\rm SO}(3,3)$ has a relative root system of type $B_{3}$ in Section 6.1. Thus, we cannot apply the results in Section 8.2. It agrees with the fact that ${\rm SL}_{4}$ is isogenous to ${\rm SO}(3,3)$, and ${\rm SL}_{4}$ does not have CAP representations, since any cuspidal representation of ${\rm SL}_{n}$ can be extended to a cuspidal representation of ${\rm GL}_{n}$ (cf. [5, p.236, line 9-10]). On the other hand, H. Narita, A. Pitale and S. Wagh [27] constructed a CAP representation on ${\rm O}(1,5)$ from Maass cusp form using the theta lift.
@@ 11.2.2. The case when $p=2$ (and $n$ is odd)
where the final identity uses a well-known formula for quadratic Gauss sums (cf. [4, Theorem 1.5.4, p.27]). ∎
We also need the following [4, p. 29]:
@@ 12. Appendix
In this section, we explicitly compute $C(S,r)=C(r,\sqrt{\beta}^{-1}\lambda)$ in (7.11) as a generalization of [30, Theorem A.1]: Let
@@
Let $I_{v}=I_{v}(r,\mu,\lambda)$. Then by the change of variables, we see that $I_{-v}=I_{v}$ and $\overline{I_{v}}=I_{v}$. So $I_{v}$ is real. Also [30, Proposition A.2] says, for any $r$,
@@
Now we explain how to modify the proof of [30, Proposition A.2] to obtain $I_{0}(2r,\mu,\lambda)=C(2r,\lambda)e^{-\mu}.$ Then together with the recurrence relation above, we obtain the theorem.
Let $s=\zeta t^{\frac{1}{2}}+\zeta^{-1}t^{-\frac{1}{2}}$, where $\zeta=e^{\sqrt{-1}\frac{\pi}{4}}$. Then $s^{2}=2+\sqrt{-1}(t-t^{-1})$ and $s\bar{s}=t+t^{-1}$. Let $\alpha=\mu+\lambda^{2}$. As in the proof of [30, Proposition A.2], we have